# Deciphering Movement: Unified Trajectory Generation Model for Multi-Agent

### v1

@@ 4.1 Benchmarks and Setup
LSTM [33] 7.10 9.02e-04 0.76 58.48 449.58 7.20 2.24e-04 0.43 34.06 228.13
Transformer [71] 6.71 2.38e-03 0.79 59.34 517.54 6.84 5.68e-04 0.42 33.01 202.10
MAT [90] 6.68 1.36e-03 0.88 58.83 483.46 6.36 4.57e-04 0.40 31.32 186.11
Naomi [43] 6.52 2.02e-03 0.81 69.10 450.66 6.77 7.66e-04 0.67 42.74 259.11
INAM [58] 6.53 3.16e-03 0.70 58.54 439.87 5.80 8.30e-04 0.39 32.10 177.04
SSSD [3] 6.18 1.82e-03 0.47 46.87 393.12 5.08 6.81e-04 0.39 23.10 122.42
GC-VRNN [85] 5.81 9.28e-04 0.37 28.08 235.99 4.95 7.12e-04 0.29 32.48 149.87
@@
Evaluation Protocol.
We generate a total of $K=20$ trajectories based on the masked trajectory input and use the minADE20 as one of the evaluation metrics. Furthermore, we include four additional metrics [90, 58] to comprehensively assess the generation quality. (1) minADE20: Calculate the minimum average displacement error between the generated trajectories and the ground truth across the 20 generated trajectories, (2) Out-of-Boundary (OOB): Measure the percentage of generated locations that fall outside the predefined court boundaries, (3) Step: Calculate the average change in step size across the generated trajectories, (4) Path-L: Calculate the average length of the trajectories for each agent, (5) Path-D: Calculate the maximum difference in trajectory lengths among the agents. Implementation details are provided in Appendix C.
4.2 Main Results
@@
LSTM [33] 186.93 4.74e-05 7.50 652.98 4542.78
Transformer [71] 170.94 6.59e-05 6.66 566.14 4269.08
MAT [90] 170.46 7.56e-05 6.45 562.44 3953.34
Naomi [43] 145.20 8.78e-05 7.47 649.62 4414.99
INAM [58] 134.86 4.04e-05 6.37 547.02 4102.37
SSSD [3] 118.71 4.51e-05 5.11 425.98 3252.66
GC-VRNN [85] 105.87 1.29e-05 4.92 506.32 3463.26
@@
Evaluation results using five metrics across three datasets are presented in Table 2 for Basketball-U and Football-U, and in Table 3 for Soccer-U. Across all datasets, our UniTraj model achieves the lowest values on the minADE20 metric, confirming its ability to generate trajectories that are closest to the ground truth. Specifically, UniTraj outperforms GC-VRNN by $17.9\%$ on the minADE20 metric for the Basketball-U dataset, $28.3\%$ for Football-U, and $10.7\%$ for Soccer-U. Regarding the OOB metric, the Mean and Median methods fill missing positions with mean and median values respectively, naturally avoiding generating out-of-bound locations. Except for these two methods, our UniTraj achieves lower rates across all three datasets, indicating its effectiveness in generating trajectories that remain within the court.
@@ 4.3 Ablation Study
We start by analyzing the contribution of each component in our UniTraj. Table 4 presents the results of the ablation study for each component. The “w/o GSM” variant omits the Ghost Spatial Masking (GSM) module, feeding only the agent embedding into the Transformer. The “w/o GSM” variant excludes the Bidirectional Temporal Scaled (BTS) module, relying solely on the bidirectional temporal mamba encoder. The “uni w/ BTS” and ‘uni w/o BTS” variants use the forward unidirectional mamba encoder, with and without the BTS module.
Comparing our full model with the “w/o GSM” variant, we see that our proposed GSM module enhances the learning of the masking pattern to boost spatial features. The comparison between the “w/o BTS” variant and the “uni w/ BTS” versus “uni w/o BTS” variants demonstrates the effectiveness of the BTS module in capturing temporal missing patterns, regardless of whether the approach is bidirectional or unidirectional. Additionally, removing the backward mamba block leads to a decrease in performance, validating the effectiveness of our bidirectional design in capturing more comprehensive temporal dependencies.
Figure 3: Ablation study results on three datasets.
@@
Figure 4: Results for different model designs.
@@
We explore how different depths of temporal mamba block affect performance. Table 4 presents the results for the Basketball-U dataset. We observe that with $L=4$, it achieves the best performance in minADE20 and Step while maintaining competitive results on other metrics. Results for other datasets are detailed in Appendix D, where similar trends are observed. Considering the balance between the number of parameters and overall performance, we set L to 4 across all three datasets.
@@ 5 Conclusion
In this work, we address the problem of modeling multi-agent trajectories by considering various situations in real practice, emphasizing the need for a general approach. To accommodate diverse real-world scenarios, we introduce a unified trajectory generation task that simultaneously handles multiple input situations. Essentially, we treat different inputs, regardless of format, as masked trajectories to generate a complete trajectory. To address various missing data scenarios, we propose a simple yet effective Ghost Spatial Masking module that integrates within the Transformer encoder and a novel Bidirectional Temporal Scaled Module embedded within the extended Bidirectional Temporal Mamba encoder. We have also curated and benchmarked three sports game datasets, Basketball-U, Football-U, and Soccer-U, to evaluate our model and establish robust baselines for future research. Extensive experiments confirm the superior performance of our approach.
@@
One limitation of our approach is the use of simple MLPs for decoding the trajectory. There is potential for improvement by developing a more powerful decoder. In addition, our sports game datasets have a fixed number of agents. How to accommodate datasets with varying numbers of agents remains an area for future exploration. We have not observed any negative societal impacts from UniTraj. Instead, we aim to advance the field by building datasets and establishing strong baselines that encourage further study into the unified trajectory generation task, with potential applications across various trajectory modeling domains.
Appendix A Datasets, Code, and Models
We provide links in the abstract to our datasets, codes, and trained checkpoints. The README.md file contains detailed instructions to download the datasets and run the codes.
Appendix B Datasets Details
@@
The base dataset is available from Stats Perform  [^3] https://www.statsperform.com/artificial-intelligence-in-sport/, which is used by recent works [90, 43, 91, 85]. The original dataset contains 104,003 training sequences and 13,464 testing sequences. Each sequence includes xy-coordinates (in feet) of the ball and 10 players (5 offensive and 5 defensive) for 8 seconds, sampled at a frequency of 6.25 Hz, resulting in sequences of length 50. The coordinates are unnormalized, with (0, 0) positioned at the bottom-left corner of the court. The court measures 94 feet in length and 50 feet in width. We first clean the dataset, removing sequences that fall outside the court boundaries. After cleaning, it consists of 93,490 training and 11,543 testing sequences. We then applied five masking strategies to generate the masks.
@@
We use the top-view scenarios from SoccerTrack dataset  [^5] https://github.com/AtomScott/SportsLabKit [63]. The dataset contains 60 tracking files, from which we use 48 for training and 12 for testing. By applying a sliding window of size 4, we extract 9,882 training sequences and 2,448 testing sequences. As the data are provided in pixel coordinates, we use the original coordinates, with the soccer field dimensions set at 3,840 by 2,160 pixels. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players.
@@
Appendix C Implementation Details and Configurations
The implementation details for different mapping functions $\varphi(\cdot)$ are provided at their first mention in the main paper. We set $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$ to balance the losses as described in Equation 11. To ensure reproducibility, we fix the random seed at 2024 across our model during training. We project the input to a dimension of 64 as in Equation 3. In our Transformer encoder, the model dimension is set to 64 and the number of heads to 8. The dimension of the latent variable $Z$ is set to 128. In our Mamba block, the state dimension is 64, the convolution kernel size is 4, the expansion value is 2, and the depth $L$ is 4. The experiments are conducted using PyTorch [56] on an NVIDIA A100 GPU. The model is trained over 100 epochs with a batch size of 128. We use the Adam optimizer [23] with an initial learning rate of 0.001, which is decayed by 0.9 every 20 epochs.
For the baseline methods Vanilla LSTM and Transformer, we adopt the same input processing as in our UniTraj. The hidden state size for LSTM and the agent embedding dimension are both set to 64. In the Transformer model, the number of heads is 8, which is the same in our UniTraj. For the MAT method [90], we use the official code available at  [^6] https://github.com/ezhan94/multiagent-programmatic-supervision. For Naomi [43], the code can be found at  [^7] https://github.com/felixykliu/NAOMI?tab=readme-ov-file. For INAM [58], we have tried our best to reproduce their methods as described in their paper, since the authors have not released the code. For SSSD [3], the code is available at  [^8] https://github.com/AI4HealthUOL/SSSD. For GC-VRNN [85], we obtained the code directly from the authors, as mentioned at  [^9] https://github.com/colorfulfuture/GC-VRNN.
@@ Appendix C Implementation Details and Configurations
Appendix D Additional Experiments

### v2

@@ 3.2 Main Results
LSTM (1997) 7.10 9.02e-04 0.76 58.48 449.58 7.20 2.24e-04 0.43 34.06 228.13
Transformer (2017) 6.71 2.38e-03 0.79 59.34 517.54 6.84 5.68e-04 0.42 33.01 202.10
MAT (2018) 6.68 1.36e-03 0.88 58.83 483.46 6.36 4.57e-04 0.40 31.32 186.11
Naomi (2019) 6.52 2.02e-03 0.81 69.10 450.66 6.77 7.66e-04 0.67 42.74 259.11
INAM (2020) 6.53 3.16e-03 0.70 58.54 439.87 5.80 8.30e-04 0.39 32.10 177.04
SSSD (2022) 6.18 1.82e-03 0.47 46.87 393.12 5.08 6.81e-04 0.39 23.10 122.42
GC-VRNN (2023b) 5.81 9.28e-04 0.37 28.08 235.99 4.95 7.12e-04 0.29 32.48 149.87
@@
LSTM (1997) 186.93 4.74e-05 7.50 652.98 4542.78
Transformer (2017) 170.94 6.59e-05 6.66 566.14 4269.08
MAT (2018) 170.46 7.56e-05 6.45 562.44 3953.34
Naomi (2019) 145.20 8.78e-05 7.47 649.62 4414.99
INAM (2020) 134.86 4.04e-05 6.37 547.02 4102.37
SSSD (2022) 118.71 4.51e-05 5.11 425.98 3252.66
GC-VRNN (2023b) 105.87 1.29e-05 4.92 506.32 3463.26
@@
Evaluation results using five metrics across three datasets are presented in Table 1 for Basketball-U and Football-U, and in Table 2 for Soccer-U. Across all datasets, our UniTraj model achieves the lowest values on the minADE20 metric, confirming its ability to generate trajectories that are closest to the ground truth. Specifically, UniTraj outperforms GC-VRNN by $17.9\%$ on the minADE20 metric for the Basketball-U dataset, $28.3\%$ for Football-U, and $10.7\%$ for Soccer-U. Regarding the OOB metric, the Mean and Median methods fill missing positions with mean and median values respectively, naturally avoiding generating out-of-bound locations. Except for these two methods, ours can achieve lower rates across all three datasets, indicating its effectiveness in generating trajectories that remain within the court.
@@ 3.3 Ablation Study
We start by analyzing the contribution of each component in our UniTraj. Table 4 presents the results of the ablation study for each component. The “w/o GSM” variant omits the Ghost Spatial Masking (GSM) module, feeding only the agent embedding into the Transformer. The “w/o BTS” excludes the Bidirectional Temporal Scaled (BTS) module, relying solely on the bidirectional temporal mamba encoder. The “uni w/ BTS” and ‘uni w/o BTS” variants use the forward unidirectional mamba encoder, with and without the BTS module.
Comparing our full model with the “w/o GSM” variant, we see that our proposed GSM module enhances the learning of the masking pattern to boost spatial features. The comparison between the “w/o BTS” variant and the “uni w/ BTS” versus “uni w/o BTS” variants demonstrates the effectiveness of the BTS module in capturing temporal missing patterns, regardless of whether the approach is bidirectional or unidirectional. Additionally, removing the backward mamba block leads to a decrease in performance, validating the effectiveness of our bidirectional design in capturing more comprehensive temporal dependencies. An interesting observation is that the contributions of certain components are different across three datasets. A potential reason is the differences in the number of players and the court sizes in these sports, which may cause the importance of spatial and temporal features to shift accordingly.
Table 3: Ablation study on three datasets.
@@
Table 4: Results for different model designs.
@@
We explore how different depths of temporal mamba block affect performance. Table 5 presents the results for the Basketball-U dataset. We observe that with $L=4$, it achieves the best performance in minADE20 and Step while maintaining competitive results on other metrics. Results for other datasets are detailed in Appendix E, where similar trends are observed. Considering the balance between the number of parameters and overall performance, we set L to 4 across all three datasets.
3.3.1 Qualitative Results
Figure 4: Qualitative comparison between advanced baselines and our method. The ball’s trajectory is shown in purple, offensive players are in green, and defensive players are in blue. Red “x” marks indicate masked locations and the starting points of the trajectories are highlighted with yellow stars.
We present one visualization example from the Basketball-U dataset, as shown in Figure 4. The trajectories generated by our method are noticeably closer to the ground truth. For the missing locations indicated by red scatter points, our method provides more accurate predictions. Furthermore, the trajectories produced by our method are significantly smoother compared to the other baselines, which highlights the effectiveness of our proposed method.
4 Related Work
Trajectory Prediction.
Trajectory prediction aims to predict the future movements of agents conditioned on their historical observations. The biggest challenge is modeling the social interactions among agents, which has driven the development of various methods. A classic approach, Social-LSTM (Alahi et al., 2016), introduced a pooling layer to enable information sharing. Subsequent methods (Vemula et al., 2018; Zhang et al., 2019; Hu et al., 2020; Xu et al., 2021) have employed similar designs to extract comprehensive interaction features. More recent studies (Kosaraju et al., 2019; Sun et al., 2020; Mohamed et al., 2020; Shi et al., 2021; Xu et al., 2022a; Li et al., 2022; Bae et al., 2022) have utilized Graph Neural Networks (GNNs), treating agents as nodes to model social interactions. The inherent uncertainty and diversity of future trajectories have also led to the adoption of generative models, including Generative Adversarial Networks (GANs) (Gupta et al., 2018; Sadeghian et al., 2019; Li et al., 2019; Amirian et al., 2019), Conditional Variational Autoencoders (CVAEs) (Mangalam et al., 2020; Xu et al., 2022b; Ivanovic & Pavone, 2019; Salzmann et al., 2020; Xu et al., 2022c), and Diffusion models (Gu et al., 2022; Mao et al., 2023; Jiang et al., 2023). However, prediction is not always the primary objective in real-world applications such as similarity analysis or action localization. Our work aims to tackle a broader range of challenges within trajectory modeling and introduces a new benchmark to better address these complexities.
Trajectory Imputation and Spatial-Temporal Recovery.
Imputation is a classic and extensively explored task, with time-series imputation receiving the most attention. Traditional statistical techniques include replacing missing values with the mean or median value (Acuna & Rodriguez, 2004), as well as methods like linear regression (Ansley & Kohn, 1984), k-nearest neighbors (Troyanskaya et al., 2001; Beretta & Santaniello, 2016), and the expectation-maximization (EM) algorithm (Ghahramani & Jordan, 1993; Nelwamondo et al., 2007). These methods often suffer from limited generalization ability due to their reliance on rigid priors. In response, more flexible frameworks have emerged, employing deep learning techniques for sequential data imputation, such as autoregressive imputation using RNNs (Yoon et al., 2018b; Cao et al., 2018), or the use of GANs, VAEs, and Diffusion models (Yoon et al., 2018a; Luo et al., 2018; 2019; Qi et al., 2020; Miao et al., 2021; Tashiro et al., 2021; Wen et al., 2024; Chen et al., 2024; Yuan & Qiao, 2024) to generate reconstructed sequences. However, only a few studies have specifically addressed trajectory imputation for multi-agent movements. For instance, AOMI (Liu et al., 2019) introduces a non-autoregressive imputation method, while GMAT (Zhan et al., 2018) develops a hierarchical model to generate macro-intent labels for sequence generation. Additionally, Graph Imputer (Omidshafiei et al., 2021) has utilized forward and backward information to model the distribution of imputed trajectories in soccer games.
The task of trajectory spatial-temporal recovery, while similar to trajectory imputation in its focus on incomplete trajectories, differs in its broader goal of reconstructing complete spatial-temporal sequences. This task requires not only filling gaps but also mining the intrinsic relations between different agents’ trajectories, which is crucial for real-world applications like arrival time estimation (Derrow-Pinion et al., 2021; Chen et al., 2022) and trajectory similarity computation (Han et al., 2021). A major benchmark involves recovering trajectory data from GPS points (Chen et al., 2011; Wei et al., 2012; 2024; Chen et al., 2023d), while few studies focus on multi-agent contexts.
Recent research has integrated trajectory imputation and prediction tasks. For instance, INAM (Qi et al., 2020) introduces an imitation learning paradigm that handles both tasks asynchronously, while GC-VRNN (Xu et al., 2023b) proposes a multi-task framework using three different GCN layers and a variational RNN to conduct trajectory imputation and prediction simultaneously. Nevertheless, these studies primarily aim to forecast trajectories based on missing observations, whereas our work aims for a more general goal that is not restricted to specific input formats or tasks. Another closely related work is Traj-MAE (Chen et al., 2023b), which introduces a continual framework to pre-train trajectory and map encoders for vehicle trajectory prediction. However, Traj-MAE focuses primarily on vehicle trajectories, which are more closely tied to maps, whereas our approach targets player trajectories in sports games, where denser social interactions occur.
State Space Models.
The core concept of state space sequence models (SSMs) (Gu et al., 2021a; b) is to link input and output sequences through a latent state. Recently, a promising variant of SSMs, Mamba architecture (Gu & Dao, 2023), which integrates time-varying parameters, has been introduced to enhance efficiency. This architecture has inspired various recent studies in different computer vision tasks. For instance, in works (Liu et al., 2024; Pei et al., 2024; Wang et al., 2024a; Behrouz et al., 2024; Zhu et al., 2024), Mamba has been applied to process image patches, learn visual representations, and achieve impressive performance in downstream tasks, proving to be an effective backbone. However, the full potential of the Mamba model in trajectory modeling remains underexplored. In this work, we extend the Mamba model to learn temporal dependencies bidirectionally and introduce a Bidirectional Temporal Scaled (BTS) module for a comprehensive scan.
@@ 5 Conclusion
In this work, we focus on the domain of sports and address the problem of modeling multi-agent trajectories by considering various situations in real practice, emphasizing the need for a general approach. To accommodate diverse real-world scenarios, we introduce a unified trajectory generation task that simultaneously handles multiple input situations. Essentially, we treat different inputs, regardless of format, as masked trajectories to generate a complete trajectory. To address various missing data scenarios, we propose a simple yet effective Ghost Spatial Masking module that integrates within the Transformer encoder and a novel Bidirectional Temporal Scaled Module embedded within the extended Bidirectional Temporal Mamba encoder. We have also curated and benchmarked three sports game datasets, Basketball-U, Football-U, and Soccer-U, to evaluate our model and establish robust baselines for future research. Extensive experiments confirm the superior performance of our approach.
@@
One limitation of our approach is the use of simple MLPs for decoding the trajectory. There is potential for improvement by developing a more powerful decoder. In addition, our sports datasets have a fixed number of agents. How to accommodate datasets with varying numbers of agents remains an area for future exploration. We have not observed any negative societal impacts from UniTraj. Instead, we hope to advance the field by building datasets and establishing strong baselines that encourage further study of the unified trajectory generation task in the sports domain.
Appendix A Appendix
Appendix B Datasets, Code, and Models
We have provided the GitHub link in the abstract to our datasets, code, and trained checkpoints. The README.md file includes detailed instructions for downloading the datasets and running the code.
Appendix C Datasets Details
@@
The base dataset is available from Stats Perform  [^3] https://www.statsperform.com/artificial-intelligence-in-sport/, which is used by recent works (Zhan et al., 2018; Liu et al., 2019; Zhan et al., 2020; Xu et al., 2023b). The original dataset contains 104,003 training sequences and 13,464 testing sequences. Each sequence includes xy-coordinates (in feet) of the ball and 10 players (5 offensive and 5 defensive) for 8 seconds, sampled at a frequency of 6.25 Hz, resulting in sequences of length 50. The coordinates are unnormalized, with (0, 0) positioned at the bottom-left corner of the court. The court measures 94 feet in length and 50 feet in width. We first clean the dataset, removing sequences that fall outside the court boundaries. After cleaning, it consists of 93,490 training and 11,543 testing sequences. We then applied five masking strategies to generate the masks.
@@
We use the top-view scenarios from SoccerTrack dataset  [^5] https://github.com/AtomScott/SportsLabKit (Scott et al., 2022). The dataset contains 60 tracking files, from which we use 48 for training and 12 for testing. By applying a sliding window of size 4, we extract 9,882 training sequences and 2,448 testing sequences. As the data are provided in pixel coordinates, we use the original coordinates, with the soccer field dimensions set at 3,840 by 2,160 pixels. Each sequence consists of trajectories for 1 ball, 11 offensive players, and 11 defensive players.
@@
Appendix D Implementation Details and Configurations
The implementation details for different mapping functions $\varphi(\cdot)$ are provided at their first mention in the main paper. We set $\lambda_{1}=\lambda_{2}=\lambda_{3}=1$ to balance the losses as described in Equation 11. To ensure reproducibility, we fix the random seed at 2024 across our model during training. We project the input to a dimension of 64 as in Equation 3. In our Transformer encoder, the model dimension is set to 64 and the number of heads to 8. The dimension of the latent variable $Z$ is set to 128. In our Mamba block, the state dimension is 64, the convolution kernel size is 4, the expansion value is 2, and the depth $L$ is 4. The experiments are conducted using PyTorch (Paszke et al., 2019) on an NVIDIA A100 GPU. The model is trained over 100 epochs with a batch size of 128. We use the Adam optimizer (Diederik & Jimmy, 2015) with an initial learning rate of 0.001, which is decayed by 0.9 every 20 epochs.
For the baseline methods Vanilla LSTM and Transformer, we adopt the same input processing as in our UniTraj. The hidden state size for LSTM and the agent embedding dimension are both set to 64. In the Transformer model, the number of heads is 8, which is the same in our UniTraj. For the MAT method (Zhan et al., 2018), we use the official code available at  [^6] https://github.com/ezhan94/multiagent-programmatic-supervision. For Naomi (Liu et al., 2019), the code can be found at  [^7] https://github.com/felixykliu/NAOMI?tab=readme-ov-file. For INAM (Qi et al., 2020), we have tried our best to reproduce their methods as described in their paper, since the authors have not released the code. For SSSD (Alcaraz & Strodthoff, 2022), the code is available at  [^8] https://github.com/AI4HealthUOL/SSSD. For GC-VRNN (Xu et al., 2023b), we obtained the code directly from the authors, as mentioned at  [^9] https://github.com/colorfulfuture/GC-VRNN.
@@ Appendix D Implementation Details and Configurations
Appendix E Additional Experiments
@@
Generalizability.
To further validate the generalizability and robustness of our proposed method, we apply it to other individual trajectory tasks. We conducted experiments on the widely used public datasets ETH-UCY Pellegrini et al. (2009); Lerner et al. (2007) and SDD (Robicquet et al., 2016) for trajectory prediction, comparing our method with recent state-of-the-art baselines: MemoNet (Xu et al., 2022b), FlowChain (Maeda & Ukita, 2023), and EqMotion (Xu et al., 2023a). As shown in Tables 10 and 12, our method outperforms FlowChain and achieves results comparable to, though slightly worse than, MemoNet and EqMotion on the ETH-UCY dataset. On the SDD dataset, our method outperforms FlowChain while performing comparably to MemoNet. However, these existing methods are challenging to adapt to other tasks, such as trajectory imputation, because their designs are specifically tailored for prediction and are not well-suited for broader applications. Additionally, our focus is on modeling players in the sports domain, which differs from the pedestrian scenarios these methods typically address.
For the imputation and recovery tasks, we conducted experiments on the recently open-sourced time-series imputation dataset Traffic-Guangzhou (Chen et al., 2018), comparing our method with the latest imputation approaches, CSBI (Chen et al., 2023c) and BayOTIDE Fang et al. (2024). The results, presented in Table 12, show that our method outperforms CSBI in both RMSE and MAE and achieves results comparable to, though slightly worse than, BayOTIDE. One possible reason for this is that time-series datasets, unlike multi-agent datasets, lack the dense, structured interactions present among sports players. Additionally, these baselines are difficult to adapt to trajectory prediction tasks. Overall, these experiments further validate the generalizability and robustness of our proposed method.
Additional Qualitative Examples
We present another visualization example from the Basketball-U dataset in Figure 5, to further demonstrate the effectiveness of our method. We can observe that the trajectories generated by our method are also closer to the ground truth and yield more accurate predictions for missing values. In addition, some baseline methods generate trajectories that are outside the basketball court boundaries, whereas our trajectories remain well-contained within the court.
Table 10: We compare our UniTraj with trajectory prediction methods on the ETH-UCY dataset. The best results are highlighted and the second best results are underlined.
Method  minADE20/minFDE20 $\downarrow$
Eth Hotel Univ Zara1 Zara2 Average
MemoNet (2022b) 0.40/0.61 0.11/0.17 0.24/0.43 0.18/0.32 0.14/0.24 0.21/0.35
FlowChain (2023) 0.55/0.99 0.20/0.35 0.29/0.54 0.22/0.40 0.20/0.34 0.29/0.52
EqMotion (2023a) 0.40/0.61 0.12/0.18 0.23/0.43 0.18/0.32 0.13/0.23 0.21/0.35
UniTraj (Ours) 0.43/0.62 0.13/0.19 0.25/0.43 0.20/0.33 0.16/0.24 0.23/0.36
Table 11: We compare our UniTraj with trajectory prediction methods on the SDD dataset. The best results are highlighted and the second best results are underlined.
Method SDD
minADE20/minFDE20 $\downarrow$
MemoNet (2022b) 8.56/12.66
FlowChain (2023) 9.93/17.17
UniTraj (Ours) 8.68/12.78
Table 12: We compare our UniTraj with time-series imputation methods on the Traffic-Guangzhou dataset. The best results are highlighted and the second best results are underlined.
Method Traffic-GuangZhou
RMSE $\downarrow$ MAE $\downarrow$
CSBI 4.790 3.182
BayOTIDE 3.820 2.687
UniTraj (Ours) 3.942 2.784
Figure 5: Qualitative comparison between advanced baselines and our method. The ball’s trajectory is shown in purple, offensive players are in green, and defensive players are in blue. Red “x” marks indicate masked locations and the starting points of the trajectories are highlighted with yellow stars.